OpenStudy (anonymous):

Integrate (9 - x^2 - x -1)dy from x = -1 to x = 2

OpenStudy (turingtest):

what problem exactly do you have with this integral?

OpenStudy (anonymous):

well the I'm actually trying to sketch a region enclosed by given curves

OpenStudy (anonymous):

but i think there is something wrong with my integration...so far I have 9y -1/3y^3 -1/2y^2 - y

OpenStudy (turingtest):

first off I assume you mean (9 - x^2 - x -1)dx

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

so then you answer is correct if you replace y's with x's

OpenStudy (anonymous):

the actual question is " Sketch the region enclosed by the given curves. Decide whether to integrate with respect to x or y. Draw a typical approximating rectangle and label its height and width. Then find the area of the region

OpenStudy (anonymous):

y = x + 1, y = 9 - x^2, x = -1, x = 2

OpenStudy (anonymous):

I am just having issues finding the area

OpenStudy (turingtest):

let's just do the integral first, yeah?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I know that I will have to take the integral [f(y) -g(y)] from c to d

OpenStudy (turingtest):

It's f(x)-g(x) I think you are having a problem keeping track of variables, remember y=f(x), so f(y) is something different and undefined here.

OpenStudy (anonymous):

I believe that is my problem, understood

OpenStudy (turingtest):

so your integral is\[\int\limits_{-1}^{2}9-x^2-x-1dx\]

OpenStudy (anonymous):

Alright, I'm so far

OpenStudy (turingtest):

your answer is right replaced with x's\[9x-{1\over3}x^3-{1\over2}x^2-x\]evaluated from -1 to 2, what do you get?

OpenStudy (anonymous):

19.2

OpenStudy (turingtest):

let me try! lol

OpenStudy (anonymous):

lol ok

OpenStudy (turingtest):

Sorry 19.5 Just the evaluation, right? As you see I hate that too, here's a site, if you don't already know it that's good for that. http://www.wolframalpha.com/ and here's your particular problem http://www.wolframalpha.com/input/?i=Integrate+%289+-+x%5E2+-+x+-1%29dx+from+-1+to+2

OpenStudy (anonymous):

Thank you so much

OpenStudy (turingtest):

wait a minute this is all wrong! I should have noticed sooner!

OpenStudy (anonymous):

No its correct I just submitted it online and it is marked correct

OpenStudy (turingtest):

did you mean it as you typed it? 9 - x^2 - x -1 are you mot missing an x? because 9-1=8 so the integration we di is wrong

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

I see what you mean, but how is it still correct

OpenStudy (turingtest):

Okay it's not ALL wrong, but the way wee did it is silly, it may be right, but we failed to notice that \[9 - x^2 - x -1=8-x^2-x\] So we integrated an extra term. It is important to see shortcuts like that, so at least look for those in the future :)

OpenStudy (anonymous):

Ok, the ah-hah moments in math...such a wonderful subject

OpenStudy (anonymous):

Thanks once again

OpenStudy (turingtest):

very welcome!